# Guided Interpretable Facial Expression Recognition via Spatial Action Unit Cues

### v1

@@ 4.3 Ablation Studies
We performed two relevant ablations, assessing how classification and localization (intepretability) performance is affected by ${\lambda}$ and at which layer the alignment loss is applied.
1- Impact of ${\lambda}$ (Fig. 5). Given different values of ${\lambda}$, we report in Fig. 5 its impact on localization (red, left curves) on CAM, and attention at layer 5 using CAM method [93] and a ResNet50 backbone [31] over the RAF-DB dataset. It is observed that increasing ${\lambda}$ improves localization at both CAM and attention. However, localization performance starts to stagnate after ${\lambda\approx 8}$. Additionally, since the alignment with action units is performed at the attention level, localization at layer-wise attention (ATT-COS) is higher than at CAM-level (CAM-COS). However, they are still both better than their corresponding vanilla cases (dashed lines). In practice, although CAMs can be used for interpretability, the attention layer is shown to be more accurate and, hence, more reliable. We also report the impact of ${\lambda}$ on classification accuracy (green, right curve). Increasing ${\lambda}$ showed to improve localization since much importance is considered for the alignment loss term. Classification accuracy also kept improving until ${\lambda\approx 4}$. However, large values of ${\lambda}$ put strong importance on localization compared to classification leading to degradation in classification accuracy. Hence, a compromise value is the one that improves localization without decreasing classification accuracy (e.g., ${\lambda\approx 4}$). Note that even large ${\lambda}$ values can still yield relatively better classification accuracy compared to the vanilla case (dashed green line).
2- At which layer should the alignment loss be applied? (Table 3). Overall, applying our action units loss term across single or multiple layers improved both classification and CAM-localization. However, its application on the last feature layer has shown to be most beneficial. Top layers often capture semantic and most discriminative regions [86]. Since action unit maps highlight discriminative ROIs, using them to build better features at top layers is a reasonable choice. In all our experiments, we use the top layer for alignment.
@@ 5 Conclusion
In this work, we have introduced an approach to make FER models more interpretable. We extended the FER model training with an additional loss that favors feature maps whose average activations are similar to facial action unit activations. As in most datasets action units localizations are not available, we estimate them by leveraging off-the-shelf facial point localizers and expert knowledge that associates action units to the corresponding emotion. By doing that, we manage to obtain a much more reliable and interpretable activation of the FER model that, in many cases, leads also to better recognition performance.
Acknowledgment
This research was supported in part by the Canadian Institutes of Health Research, the Natural Sciences and Engineering Research Council of Canada, and the Digital Research Alliance of Canada (alliancecan.ca).

### v3

@@ 4.3 Ablation Studies
We performed two relevant ablations, assessing how classification and localization (interpretability) performance is affected by ${\lambda}$ and at which layer the alignment loss is applied.
1- Impact of ${\lambda}$ (Fig. 5). Given different values of ${\lambda}$, we report in Fig. 5 its impact on localization (red, left curves) on CAM, and attention at layer 5 using CAM method [93] and a ResNet50 backbone [31] over the RAF-DB dataset. It is observed that increasing ${\lambda}$ improves localization at both CAM and attention. However, localization performance starts to stagnate after ${\lambda\approx 8}$. Additionally, since the alignment with AUs is performed at the attention level, localization at layer-wise attention (ATT-COS) is higher than at CAM-level (CAM-COS). However, they are still better than their corresponding vanilla cases (dashed lines). In practice, although CAMs can be used for interpretability, the attention layer is more accurate and, hence, more reliable. We also report the impact of ${\lambda}$ on classification accuracy (green, right curve). Increasing ${\lambda}$ showed to improve localization since much importance is considered for the alignment loss term. Classification accuracy also kept improving until ${\lambda\approx 4}$. However, large values of ${\lambda}$ put substantial importance on localization compared to classification, leading to degradation in classification accuracy. Hence, a compromise value is the one that improves localization without decreasing classification accuracy (e.g., ${\lambda\approx 4}$). Even large ${\lambda}$ values can yield relatively better classification accuracy than the vanilla case (dashed green line).
2- Layer where alignment loss is applied (Table 3). Overall, applying our AU loss term across single or multiple layers improved both classification and CAM-localization. However, its application on the last feature layer has shown to be most beneficial. Top layers often capture semantic and most discriminative regions [86]. Since AU maps highlight discriminative ROIs, using them to build better features at top layers is a reasonable choice. In all our experiments, we use the top layer for alignment.
@@ 5 Conclusion
In this work, we have introduced an approach to make FER models more interpretable. We extended the FER model training with an additional loss that favors feature maps whose average activations are similar to facial AU activations. As in most datasets AU localizations are not available, we estimate them by leveraging off-the-shelf facial point localizers and expert knowledge that associates AUs to the corresponding emotion. By doing that, we manage to obtain a much more reliable and interpretable activation of the FER model that, in many cases, also leads to better recognition performance.
Acknowledgments
This work was supported in part by the Fonds de recherche du Québec – Santé (FRQS), the Natural Sciences and Engineering Research Council of Canada (NSERC), Canada Foundation for Innovation (CFI), and the Digital Research Alliance of Canada.